# Continuous-time q-learning in jump-diffusion models under Tsallis entropy

### v3

@@ 5.2 A non-LQ optimal repo rate control problem
${\rm s.t.}\leavevmode\nobreak\ X_{s}^{\pi}$ $=x+\int_{t}^{s}\int_{\mathds{R}^{2}}(u_{1}\mu_{1}+u_{2}\mu_{2})X_{v}^{\pi}\pi_{v}(du)dv+\int_{t}^{s}\sigma X_{v}^{\pi}dW_{v}$ (5.15)
$\quad-\int_{0}^{t}\nu X_{v-}^{\pi}dN_{v},\leavevmode\nobreak\ \forall s\in[t,T].$
@@
$\int_{\mathds{R}^{2}}u_{1}^{k}\widehat{\pi}(u|t,x)du=\begin{cases}\displaystyle\frac{\mu_{1}}{2Ax^{h}}\alpha^{*}(t),\quad k=1,\\[11.99998pt] \displaystyle\frac{(p-1)\tilde{\psi}(t,x)}{2A(2p-1)x^{2h}}+\left(\frac{\mu_{1}}{2Ax^{h}}\alpha^{*}(t)\right)^{2},\quad k=2,\end{cases}$
@@
$\int_{\mathds{R}^{2}}u_{2}^{k}\widehat{\pi}(u|t,x)du=\begin{cases}\displaystyle\frac{\mu_{2}}{2Bx^{h}}\alpha^{*}(t),\quad k=1,\\[11.99998pt] \displaystyle\frac{(p-1)\tilde{\psi}(t,x)}{2B(2p-1)x^{2h}}+\left(\frac{\mu_{2}}{2Bx^{h}}\alpha^{*}(t)\right)^{2},\quad k=2.\end{cases}$
@@
$\int_{\mathds{R}^{2}}\frac{1}{p-1}\left(\widehat{\pi}(u|t,x)-\widehat{\pi}(u|t,x)^{p}\right)du=\frac{1}{p-1}-\int_{\mathds{R}^{2}}\frac{\widehat{\pi}(u|t,x)^{p}}{p-1}du=\frac{1}{p-1}-\frac{\tilde{\psi}(t,x)}{\gamma(2p-1)}.$ (5.19)
@@
$\begin{cases}\displaystyle\frac{(\alpha^{*}(t))^{\prime}}{h}+\left(\frac{\sigma^{2}}{2}(h-1)+\lambda\frac{(1-\nu)^{h}-1}{h}\right)\alpha^{*}(t)-\frac{4}{3}\sqrt{\frac{\gamma}{\pi}}{(AB)}^{\frac{1}{4}}=0,\leavevmode\nobreak\ \alpha^{*}(T)=1,\\[8.00003pt] \displaystyle(\beta^{*}(t))^{\prime}+\left(\frac{\mu_{1}^{2}}{4A}+\frac{\mu_{2}^{2}}{4B}\right)(\alpha^{*}(t))^{2}+\gamma=0,\leavevmode\nobreak\ \beta^{*}(T)=0.\end{cases}$
@@
Notably, the explicit results in Proposition 5.5 are exclusive to the Tsallis entropy when $p=2$, while no exact parameterization is available under the conventional Shannon entropy in this example. The optimal policy given by (5.22) is also a two-dimensional q-Gaussian distribution with a compact support set (see Figure 5). In fact, for $(t,x)\in[0,T]\times(0,+\infty)$, we have
@@
Figure 5: The optimal policy $(u_{1},u_{2})\to\widehat{\pi}(u_{1},u_{2})$. The model parameters are set to be $\lambda=1,\leavevmode\nobreak\ \sigma=1,\leavevmode\nobreak\ \nu=0.5,\leavevmode\nobreak\ A=B=1,\leavevmode\nobreak\ \mu_{1}=\mu_{2}=0.5,\leavevmode\nobreak\ h=1.5,\leavevmode\nobreak\ \gamma=1,\leavevmode\nobreak\ t=1,\leavevmode\nobreak\ T=2,\leavevmode\nobreak\ x=1$.
@@
This implies the convergence of the borrowing and lending policy to a constant strategy that $(\xi_{t},\eta_{t})\xrightarrow[\gamma\to 0]{L^{2}}\left(\frac{\mu_{1}}{2Ax^{h}}\alpha^{*}(t),\frac{\mu_{2}}{2Bx^{h}}\alpha^{*}(t)\right)$ for $(t,x)\in[0,T]\times\mathds{R}_{+}$.
@@
Simulator: With the above parameterized value function and q-function, we next apply the Algorithm 1. We use the acceptance-rejection sampling method to generate the control pair $(u^{1}_{t},u^{2}_{t})$ from the q-Gaussian distribution with density function given by (5.12) at time $t$. Then the control pair $(u^{1}_{t},u^{2}_{t})$ is used to generate sample trajectories through the following simulator
@@
$\alpha_{\theta_{1}}(k)$ $=\frac{0.0023}{{\rm linspace}_{(1,90,N)}(k)},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for }1\leq k\leq N,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{\theta_{2}}(k)=\frac{0.0325}{{\rm linspace}_{(1,90,N)}(k)},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for }1\leq k\leq N,$
$\alpha_{\theta_{3}}(k)$ $=\frac{0.0017}{{\rm linspace}_{(1,60,N)}(k)},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for }1\leq k\leq N,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{\zeta_{1}}(k)=\frac{0.0026}{{\rm linspace}_{(1,50,N)}(k)},\leavevmode\nobreak\ \text{for }1\leq k\leq N,$
$\alpha_{\zeta_{2}}(k)$ $=\begin{cases}\displaystyle 0.005,\leavevmode\nobreak\ \text{if }1\leq k\leq 5200,\\ \displaystyle\frac{0.01}{{\rm linspace}_{(1,500,N)}(k)},\leavevmode\nobreak\ \text{if }5200<k\leq N,\end{cases}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{\zeta_{3}}(k)=\begin{cases}\displaystyle 0.002,\leavevmode\nobreak\ \text{if }1\leq k\leq 6100,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,500,N)}(k)},\leavevmode\nobreak\ \text{if }6100<k\leq N,\end{cases}$
$\alpha_{\zeta_{4}}(k)$ $=\frac{0.0046}{{\rm linspace}_{(1,150,N)}(k)},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for }1\leq k\leq N,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{\zeta_{5}}(k)=\frac{0.0045}{{\rm linspace}_{(1,150,N)}(k)},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for }1\leq k\leq N,$
$\alpha_{\zeta_{6}}(k)$ $=\begin{cases}\displaystyle\frac{0.015}{{\rm linspace}_{(1,80,N)}(k)},\leavevmode\nobreak\ \text{if }1\leq k\leq 8000,\\ \displaystyle 0.00001,\leavevmode\nobreak\ \text{if }8000<k\leq N,\end{cases}$
Table 2: Parameters used in the simulator.
@@
$\alpha_{\chi_{1}}(k)$ $=\frac{0.026}{{\rm linspace_{(1,100,N)}}(k)},\leavevmode\nobreak\ \text{for }1\leq k\leq N,\leavevmode\nobreak\ \alpha_{\chi_{2}}(k)=\frac{0.05}{{\rm linspace_{(1,500,N)}}(k)},\leavevmode\nobreak\ \text{for }1\leq k\leq N,$
$\alpha_{\chi_{3}}(k)$ $=\begin{cases}\displaystyle 0.002,\leavevmode\nobreak\ \text{if }1\leq k\leq 6100,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,500,N)}(k)},\leavevmode\nobreak\ \text{if }6100\leq k\leq N,\end{cases}\leavevmode\nobreak\ \alpha_{\chi_{4}}(k)=\frac{0.00461}{{\rm linspace_{(1,150,N)}}(k)},\leavevmode\nobreak\ \text{for }1\leq k\leq N,$
$\alpha_{\chi_{5}}(k)$ $=\frac{0.005}{{\rm linspace_{(1,200,N)}}(k)},\leavevmode\nobreak\ \text{for }1\leq k\leq N,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{\chi_{6}}(k)=\begin{cases}\displaystyle\frac{0.0015}{{\rm linspace}_{(1,80,N)}(k)},\leavevmode\nobreak\ \text{if }1\leq k\leq 8000,\\ \displaystyle 0.00001,\leavevmode\nobreak\ \text{if }80001\leq k\leq N.\end{cases}$

### v4

@@ 5.2 A non-LQ optimal repo rate control problem
${\rm s.t.}\penalty 10000\ X_{s}^{\pi}$ $=x+\int_{t}^{s}\int_{\mathds{R}^{2}}(u_{1}\mu_{1}+u_{2}\mu_{2})X_{v}^{\pi}\pi_{v}(du)dv+\int_{t}^{s}\sigma X_{v}^{\pi}dW_{v}$ (5.17)
$\quad-\int_{0}^{t}\nu X_{v-}^{\pi}dN_{v},\penalty 10000\ \forall s\in[t,T].$
@@
$\int_{\mathds{R}^{2}}u_{1}^{k}\widehat{\pi}(u|t,x)du=\begin{cases}\displaystyle\frac{\mu_{1}}{2Ax^{h}}\alpha^{*}(t),\quad k=1,\\[11.99998pt] \displaystyle\frac{(p-1)\tilde{\psi}(t,x)}{2A(2p-1)x^{2h}}+\left(\frac{\mu_{1}}{2Ax^{h}}\alpha^{*}(t)\right)^{2},\quad k=2,\end{cases}$ (5.21)
@@
$\int_{\mathds{R}^{2}}u_{2}^{k}\widehat{\pi}(u|t,x)du=\begin{cases}\displaystyle\frac{\mu_{2}}{2Bx^{h}}\alpha^{*}(t),\quad k=1,\\[11.99998pt] \displaystyle\frac{(p-1)\tilde{\psi}(t,x)}{2B(2p-1)x^{2h}}+\left(\frac{\mu_{2}}{2Bx^{h}}\alpha^{*}(t)\right)^{2},\quad k=2.\end{cases}$ (5.22)
@@
$\int_{\mathds{R}^{2}}\frac{1}{p-1}\left(\widehat{\pi}(u|t,x)-\widehat{\pi}(u|t,x)^{p}\right)du=\frac{1}{p-1}-\frac{\tilde{\psi}(t,x)}{\gamma(2p-1)}.$ (5.23)
@@
$\begin{cases}\displaystyle\frac{(\alpha^{*}(t))^{\prime}}{h}+\left(\frac{\sigma^{2}}{2}(h-1)+\lambda\frac{(1-\nu)^{h}-1}{h}\right)\alpha^{*}(t)-\frac{4}{3}\sqrt{\frac{\gamma}{\pi}}{(AB)}^{\frac{1}{4}}=0,\penalty 10000\ \alpha^{*}(T)=1,\\[8.00003pt] \displaystyle(\beta^{*}(t))^{\prime}+\left(\frac{\mu_{1}^{2}}{4A}+\frac{\mu_{2}^{2}}{4B}\right)(\alpha^{*}(t))^{2}+\gamma=0,\penalty 10000\ \beta^{*}(T)=0.\end{cases}$
@@
Notably, the explicit results in Proposition 5.5 are exclusive to the Tsallis entropy when $p=2$, while no exact parameterization is available under the conventional Shannon entropy in this example. The optimal policy given by (5.26) is also a two-dimensional $p$-Gaussian distribution with a compact support set (see Figure 5). The mean and variance of the policy $(u_{1},u_{2})$ are given by
${\rm mean}(u_{1},u_{2})=\left(\int_{\mathds{R}^{2}}u_{1}\widehat{\pi}(u)du,\int_{\mathds{R}^{2}}u_{2}\widehat{\pi}(u)du\right)=\frac{\alpha^{*}(t)}{2x^{h}}\left(\frac{\mu_{1}}{A},\frac{\mu_{2}}{B}\right),$ (5.27)
${\rm Var}(u_{1},u_{2})=\left(\int_{\mathds{R}^{2}}u_{1}^{2}\widehat{\pi}(u)du-\left(\int_{\mathds{R}^{2}}u_{1}\widehat{\pi}(u)du\right)^{2},\int_{\mathds{R}^{2}}u_{2}^{2}\widehat{\pi}(u)du-\left(\int_{\mathds{R}^{2}}u_{2}\widehat{\pi}(u)du\right)^{2}\right)$
$\qquad\qquad=\frac{(p-1)\tilde{\psi}(t,x)}{2(2p-1)x^{2h}}\left(\frac{1}{A},\frac{1}{B}\right).$ (5.28)
In fact, for $(t,x)\in[0,T]\times(0,+\infty)$, we have
@@
Figure 5: The optimal policy $(u_{1},u_{2})\to\widehat{\pi}(u_{1},u_{2})$. The model parameters are set to be $\lambda=1,\penalty 10000\ \sigma=1,\penalty 10000\ \nu=0.5,\penalty 10000\ A=B=1,\penalty 10000\ \mu_{1}=\mu_{2}=0.5,\penalty 10000\ h=1.5,\penalty 10000\ \gamma=1,\penalty 10000\ t=1,\penalty 10000\ T=2,\penalty 10000\ x=1$.
@@
This implies the convergence of the borrowing and lending policy to a deterministic strategy that $(\xi_{t},\eta_{t})\xrightarrow[\gamma\to 0]{L^{2}}\left(\frac{\mu_{1}}{2Ax^{h}}\alpha^{*}(t),\frac{\mu_{2}}{2Bx^{h}}\alpha^{*}(t)\right)$ for $(t,x)\in[0,T]\times\mathds{R}_{+}$, which is the optimal strategy of the classical stochastic control problem.
@@
Simulator: With the above parameterized value function and q-function, we next apply the Algorithm 1. We use the acceptance-rejection sampling method to generate the control pair $(u^{1}_{t},u^{2}_{t})$ from the $p$-Gaussian distribution with density function given by (5.14) at time $t$. Then the control pair $(u^{1}_{t},u^{2}_{t})$ is used to generate sample trajectories through the following simulator
@@
$\alpha_{\theta_{1}}(k)$ $=\frac{0.0023}{{\rm linspace}_{(1,90,N)}(k)},\penalty 10000\ \penalty 10000\ \text{for }1\leq k\leq N,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \alpha_{\theta_{2}}(k)=\frac{0.0325}{{\rm linspace}_{(1,90,N)}(k)},\penalty 10000\ \penalty 10000\ \text{for }1\leq k\leq N,$
$\alpha_{\theta_{3}}(k)$ $=\frac{0.0017}{{\rm linspace}_{(1,60,N)}(k)},\penalty 10000\ \penalty 10000\ \text{for }1\leq k\leq N,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \alpha_{\zeta_{1}}(k)=\frac{0.0026}{{\rm linspace}_{(1,50,N)}(k)},\penalty 10000\ \text{for }1\leq k\leq N,$
$\alpha_{\zeta_{2}}(k)$ $=\begin{cases}\displaystyle 0.005,\penalty 10000\ \text{if }1\leq k\leq 5200,\\ \displaystyle\frac{0.01}{{\rm linspace}_{(1,500,N)}(k)},\penalty 10000\ \text{if }5200<k\leq N,\end{cases}\penalty 10000\ \penalty 10000\ \penalty 10000\ \alpha_{\zeta_{3}}(k)=\begin{cases}\displaystyle 0.002,\penalty 10000\ \text{if }1\leq k\leq 6100,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,500,N)}(k)},\penalty 10000\ \text{if }6100<k\leq N,\end{cases}$
$\alpha_{\zeta_{4}}(k)$ $=\frac{0.0046}{{\rm linspace}_{(1,150,N)}(k)},\penalty 10000\ \penalty 10000\ \text{for }1\leq k\leq N,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \alpha_{\zeta_{5}}(k)=\frac{0.0045}{{\rm linspace}_{(1,150,N)}(k)},\penalty 10000\ \penalty 10000\ \text{for }1\leq k\leq N,$
$\alpha_{\zeta_{6}}(k)$ $=\begin{cases}\displaystyle\frac{0.015}{{\rm linspace}_{(1,80,N)}(k)},\penalty 10000\ \text{if }1\leq k\leq 8000,\\ \displaystyle 0.00001,\penalty 10000\ \text{if }8000<k\leq N,\end{cases}$
Table 2: Parameters
@@
$\alpha_{\chi_{1}}(k)$ $=\frac{0.026}{{\rm linspace_{(1,100,N)}}(k)},\penalty 10000\ \text{for }1\leq k\leq N,\penalty 10000\ \alpha_{\chi_{2}}(k)=\frac{0.05}{{\rm linspace_{(1,500,N)}}(k)},\penalty 10000\ \text{for }1\leq k\leq N,$
$\alpha_{\chi_{3}}(k)$ $=\begin{cases}\displaystyle 0.002,\penalty 10000\ \text{if }1\leq k\leq 6100,\\ \displaystyle\frac{0.005}{{\rm linspace}_{(1,500,N)}(k)},\penalty 10000\ \text{if }6100\leq k\leq N,\end{cases}\penalty 10000\ \alpha_{\chi_{4}}(k)=\frac{0.00461}{{\rm linspace_{(1,150,N)}}(k)},\penalty 10000\ \text{for }1\leq k\leq N,$
$\alpha_{\chi_{5}}(k)$ $=\frac{0.005}{{\rm linspace_{(1,200,N)}}(k)},\penalty 10000\ \text{for }1\leq k\leq N,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \alpha_{\chi_{6}}(k)=\begin{cases}\displaystyle\frac{0.0015}{{\rm linspace}_{(1,80,N)}(k)},\penalty 10000\ \text{if }1\leq k\leq 8000,\\ \displaystyle 0.00001,\penalty 10000\ \text{if }80001\leq k\leq N.\end{cases}$
@@
6 Conclusions
This paper studies the continuous-time q-learning in jump-diffusion models under Tsallis entropy regularization. By considering the general entropy beyond the Shannon type, we obtain optimal policies that are not necessarily Gibbs measures but with compact support. We establish the martingale characterization of the q-function and devise several q-learning algorithms. The effectiveness of our algorithms is demonstrated in two financial applications where the parameter iterations exhibit desirable convergence. One appealing future study is to investigate the continuous-time RL algorithms under Tsallis entropy for mean field game and mean field control problems. Moreover, as the sampling for the resulting optimal policy with compact support might be challenging in high-dimensional scenarios, another interesting future research direction is to examine some tailor-made sampling methods, such as Neal 2003, Tong et al. 2020, in our RL algorithms for some high-dimensional financial applications.